OpenStudy (anonymous):

3. Ammonia, NH3, is a typical ingredient in household cleaners. It is produced through a combination reaction involving N2(g) and H2(g). If 24.0 mol of H2(g) react with excess N2(g), how many moles of ammonia are produced?

OpenStudy (anonymous):

the answer is 16 moles

OpenStudy (anonymous):

i know i need the method

OpenStudy (anonymous):

see the balanced equation is N2 + 3(H2) = 2(NH3) SO, 3 moles H2 give 2 moles NH3 then 1 mole H2 gives 2/3 moles NH3 then 24 moles H2 gives 2/3*24=16 moles NH3

OpenStudy (anonymous):

formula

OpenStudy (anonymous):

There is no formula it is just unitary method!!

OpenStudy (anonymous):

R u satisfied???

OpenStudy (anonymous):

8. Explain how the atoms of one isotope differ from those of other isotopes of the same element. o Isotopes differs because of the number of neutrons in the nucleus. o Some combinations of protons and neutrons produce an unstable nucleus. o They may also differ by being radioactive.

OpenStudy (anonymous):

the answer is option 1

OpenStudy (anonymous):

no there are no options

OpenStudy (anonymous):

it was answer for my questions if it is wrong please correct it and i also have to dimensional analysiis with it

OpenStudy (anonymous):

hmmmm see, no. of protons in any isotope of same element remains same, so i dont think ur 2nd point is apt bt 1st n 3rd point are ok

OpenStudy (anonymous):

can you tell me the dimensional analysis for it

OpenStudy (anonymous):

What do u mean by dimensional analysis???

OpenStudy (anonymous):

my teacher says dimensional analysis is needed

OpenStudy (anonymous):

u want dimensional analysis of NH3 ques????

OpenStudy (anonymous):

no isotopes one

OpenStudy (preetha):

There is no dimensional analysis for the isotope q.

OpenStudy (anonymous):

Yup i thought the same @Preetha

OpenStudy (preetha):

And lily, you should give Lord a medal for his help on the first question

OpenStudy (anonymous):

that hat iam saying to her she is not ready to listen

OpenStudy (anonymous):

thnxx again @Preetha

OpenStudy (preetha):

sorry, I cant think of how to use dimensional analysis.

OpenStudy (anonymous):

10. Convert 4.3 kilograms to centigrams using dimensional analysis. can you tell me the answer for this one with the formula

OpenStudy (anonymous):

k = kilo c = centi g= gram 1 kg = 1000 g : 1 g = 100 cg so, 4.3 kg = 4300 g = 4300(100 cg) = 4.3 x 10^5 cg.

OpenStudy (anonymous):

i gave her this answer she says according to dimensional analyisis

OpenStudy (anonymous):

Ur answer and method is correct but sorry i dont know dimensional analysis method

OpenStudy (mertsj):

\[4.3 kg \times \frac{1000g}{1 kg}\times \frac{100cg}{1 g}=4.3 \times10^5 cg\]